# Basic representations of genus zero nonabelian Hodge spaces

### v2

@@ 4.2. Painlevé II
The standard rank 2 Painlevé II Lax representation has only one irregular singularity, at infinity, given by a pole of order 4.
The corresponding fission forest is
111122223333
It admits representatives given by irregular classes of the form:
@@
$\bm{\Theta}^{II}=\bm{\breve{\Theta}}^{II}=\langle\alpha z^{3}\rangle_{\infty}+\langle\beta z^{3}\rangle_{\infty},$
with $\alpha\neq\beta$.
To have a minimal rank of the generic class of representations and recover the other known Lax representations for Painlevé II, we need to assume that $\beta=0$, to which we can always reduce up to tensoring with a rank one local system (but beware this twist is by an exponential factor of slope 3, so does not belong to the $\operatorname{SL}_{2}(\mathbb{C})$ action. However, since it is at infinity, it preserves the diagram). Let us thus consider an irregular class of the form
$\bm{\breve{\Theta}}^{II}_{std}=\langle\alpha z^{3}\rangle_{\infty}+\langle 0\rangle_{\infty},$
with $\alpha\neq 0$ and study the classes of its nearby representations.
The generic class of representations has rank 3, and admits representatives of the form
$\bm{\breve{\Theta}}^{II}_{gen}=\langle\lambda z^{2}+\alpha z^{3/2}\rangle_{\infty}+\langle\mu z^{2}\rangle_{\infty},$
with $\lambda\neq\mu$, $\alpha\neq 0$.
The generic fission tree is
111122223/2323/23 / 2
We have $k=2$ Fourier sphere coefficients, $N_{1}=N_{1}^{+}=\{\langle\lambda z^{2}+\alpha z^{3/2}\rangle_{\infty}\}$, $N_{2}=N_{2}^{-}=\{\langle\mu z^{2}\rangle_{\infty}\}$. The two principal subtrees are not isomorphic, so the two nongeneric classes of representations have distinct forests.
The standard representation corresponds to the nongeneric class of representations for $i=1$. The second nongeneric class of representations also has rank 2, and admits representatives of the form
$\bm{\breve{\Theta}}^{II}_{2}=\langle\alpha z^{3/2}\rangle_{\infty}+\langle 0\rangle_{0}.$
We notice that it corresponds precisely the other known Lax representation for Painlevé II due to Flaschka-Newell [25], also sometimes referred to as degenerate Painlevé IV.
Finally the diagram is
and the basic representations correspond to different readings of the diagram, as indicated on the figure below:
$\langle\alpha z^{3}\rangle_{\infty}$$\langle 0\rangle_{\infty}$$\bm{\breve{\Theta}}^{II}_{std}$$\langle\lambda z^{2}+\alpha z^{3/2}\rangle_{\infty}$$\langle\mu z^{2}\rangle_{\infty}$$\bm{\breve{\Theta}}^{II}_{gen}$$\langle\alpha z^{3/2}\rangle_{\infty}$$\langle 0\rangle_{0}$$\bm{\breve{\Theta}}^{II}_{2}$
4.3. Painlevé III
The standard rank 2 Painlevé III Lax representation has two irregular singularities, both corresponding to a second order pole, one of them being at infinity.
The fission forest is
$1$
The general form of a global irregular class with this fission forest and a singularity at infinity is  [^1] We will denote Painlevé III, degenerate Painlevé III and doubly degenerate Painlevé III by $III(2)$, $III(1)$ and $III(0)$ respectively, the numbers 2,1,0 referring to the number of parameters of each equation.
$\bm{\Theta}^{III(2)}_{std}=\langle\alpha z\rangle_{\infty}+\langle\beta z\rangle_{\infty}+\langle\alpha^{\prime}z^{-1}_{a}\rangle_{a}+\langle\beta^{\prime}z_{a}^{-1}\rangle_{a},$
with $\alpha,\beta,\alpha^{\prime},\beta^{\prime}\in\mathbb{C}$, $a\neq\infty$, $\alpha\neq\beta$, $\alpha^{\prime}\neq\beta^{\prime}$.
If $(E,\nabla)$ is a connection of this type, we can always apply a twist by a rank one local system to set $\beta^{\prime}$ to zero, to obtain a connection with a modified global irregular class of the form (whose fission forest is the one drawn on Fig. LABEL:fig:PIII_tree_and_diagram)
$\bm{\breve{\Theta}}^{III(2)}_{std}=\langle\alpha z\rangle_{\infty}+\langle\beta z\rangle_{\infty}+\langle\alpha^{\prime}z_{a}^{-1}\rangle_{a}.$
with $\alpha^{\prime}\neq 0$.
Let us study the classes of nearby representations of $\bm{\breve{\Theta}}^{III(2)}_{std}$. The generic class of representations has rank 4, and admits representatives of the following form
$\bm{\breve{\Theta}}^{III(2)}_{gen}=\langle\lambda z^{2}+\alpha z\rangle_{\infty}+\langle\lambda z^{2}+\beta z\rangle_{\infty}+\langle\mu z^{2}+\gamma z^{1/2}\rangle_{\infty},$
with $\lambda\neq\mu$, $\alpha\neq\beta$, $\gamma\neq 0$.
We have $k=2$ Fourier sphere coefficients, $N_{1}=N_{1}^{-}=\{\langle\lambda z^{2}+\alpha z\rangle_{\infty},\langle\lambda z^{2}+\beta z\rangle_{\infty}\}$, $N_{2}=N_{2}^{-}=\{\langle\mu z^{2}+\gamma z^{1/2}\rangle_{\infty}\}$.
The corresponding fission tree is
111122221/2121/21 / 2
Since the two principal subtrees are not isomorphic, the forests of the two corresponding nongeneric representations do not coincide. The standard representation is the nongeneric representation for $i=1$. The other nongeneric class of representations has rank 2 and corresponds to irregular classes of the form
$\bm{\breve{\Theta}}^{III(2)}_{2}=\langle\alpha z^{1/2}\rangle_{\infty}+\langle 0\rangle_{a}+\langle 0\rangle_{b},$
with $a\neq b$, $\alpha\neq 0$. It is none other but the Flaschka-Newell Lax representation for Painlevé III [25], also sometimes referred to as degenerate Painlevé V.
Finally, the diagram is
and the classes of basic representations correspond to different readings, as represented on the figure below
@@ 4.3. Painlevé III
$\langle\lambda z^{2}+\alpha z\rangle_{\infty}$$\langle\mu z^{2}+\gamma z^{1/2}\rangle_{\infty}$$\langle\lambda z^{2}+\beta z\rangle_{\infty}$$\bm{\breve{\Theta}}^{III(2)}_{gen}$$\langle\alpha z\rangle_{\infty}$$\langle\alpha^{\prime}z_{a}\rangle_{a}$$\langle\beta z\rangle_{\infty}$$\bm{\breve{\Theta}}^{III(2)}_{std}$$\langle 0\rangle_{a}$$\langle\alpha z^{1/2}\rangle_{\infty}$$\langle 0\rangle_{b}$$\bm{\breve{\Theta}}^{III(2)}_{2}$
4.4. Degenerate Painlevé III
The standard Lax representation for degenerate Painlevé III, as for nondegenerate Painlevé III, has two irregular singularities, one of them at infinity, both corresponding to poles of order 2, but with a nilpotent leading term for the singularity at infinity, giving rise in the Turritin-Levelt normal form to a single Stokes circle of slope $1/2$ [40].
The general form of a global irregular class of this type is
@@ 4.4. Degenerate Painlevé III
$\bm{\Theta}^{III(1)}_{std}=\langle\beta z^{1/2}\rangle_{\infty}+\langle\alpha z_{a}^{-1}\rangle_{a}+\langle\alpha^{\prime}z_{a}^{-1}\rangle_{a},$
with $a\neq\infty$, $\beta\neq 0$, and $\alpha\neq\alpha^{\prime}$.
The corresponding fission forest is
11111/2121/21 / 2
Given a connection of this type, we can always apply a twist by a rank one local system to set $\alpha^{\prime}=0$ and kill the formal monodromy of $\langle\alpha^{\prime}z_{a}^{-1}\rangle_{a}$, and obtain a connection with a global modified irregular class of the form:
$\bm{\breve{\Theta}}^{III(1)}_{std}=\langle\beta z^{1/2}\rangle_{\infty}+\langle\alpha z^{-1}\rangle_{a},$
with $\beta\neq 0$, $\alpha\neq 0$.
Now let us consider the classes of nearby representations of $\bm{\breve{\Theta}}^{III(1)}_{std}$. The generic class of representations has rank 4, and admits representatives of the form
$\bm{\breve{\Theta}}^{III(1)}_{gen}=\langle\lambda z^{2}+\alpha z^{1/2}\rangle_{\infty}+\langle\mu z^{2}+\beta z^{1/2}\rangle_{\infty}$
with $\lambda\neq\mu$, $\alpha\neq 0$ and $\beta\neq 0$.
The generic fission tree is
111122221/2121/21 / 2
We have $k=2$ Fourier sphere coefficients, $N_{1}=N_{1}^{-}=\{\langle\lambda z^{2}+\alpha z^{1/2}\rangle_{\infty}\}$, and $N_{2}=N_{2}^{-}=\{\langle\mu z^{2}+\beta z^{1/2}\rangle_{\infty}\}$.
@@
Furthermore, the two principal subtrees are isomorphic, so the two nongeneric classes of representations have the same forest, and there are only two distinct forests, the generic one and the nongeneric one.
Finally, the diagram is
444411-1- $-1$
@@
and the distinct forests correspond to different types of readings of the diagram, as indicated on the figure below:

### v3

@@ 4.3.1 Standard representation
The standard rank 2 Painlevé II Lax representation has one irregular untwisted singularity corresponding to a pole of order 4.
This corresponds to the following fission forest $\mathbf{F}^{\rm II}_{\rm std}$:
112233
The irregular classes with fission forest $\mathbf{F}^{\rm II}_{\rm std}$ and minimal Poincaré–Katz rank (equal to 3) are all irregular classes of the form
$\bm{\Theta}_{\rm std}^{\rm II}=\bigl\langle\alpha_{3}z_{a}^{-3}+\alpha_{2}z_{a}^{-2}+\alpha_{1}z_{a}^{-1}\bigr\rangle_{a}+\bigl\langle\beta_{3}z_{a}^{-3}+\beta_{2}z_{a}^{-2}+\beta_{1}z_{a}^{-1}\bigr\rangle_{a},$
with $a\in\mathbb{P}^{1}$ and $\alpha_{3}\neq\beta_{3}$.
Given a connection on $\mathbb{P}^{1}$ with such an irregular class, the rank of its generic class of nearby representations is minimal when $a=\infty$ and (say) $\beta_{3}=0$ (to which we can always reduce up to applying a Möbius transformation and tensoring with a rank one connection).
Let us thus consider a (modified) irregular class of the form
@@
$\bm{\breve{\Theta}}^{\rm II}_{\rm std}=\bm{\Theta}^{\rm II}_{\rm std}=\bigl\langle\alpha_{3}z^{3}+\alpha_{2}z^{2}+\alpha_{1}z\bigr\rangle_{\infty}+\bigl\langle\beta_{2}z^{2}+\beta_{1}z\bigr\rangle_{\infty},$
with $\alpha_{3}\neq 0$ and determine the classes of its nearby representations.
4.3.2 Generic representation
The short fission tree $\mathbf{T}^{\rm II}$ of $\bm{\breve{\Theta}}^{\rm II}_{\rm std}$ is
11223/23/21/21/2
An irregular class of generic form has fission tree $\mathbf{T}^{\rm II}$ if and only if it is of the form
@@ 4.3.2 Generic representation
$\bm{\breve{\Theta}}^{\rm II}_{\rm gen}=\bigl\langle\lambda^{\prime}z^{2}+\alpha^{\prime}_{3/2}z^{3/2}+\alpha^{\prime}_{1}z+\alpha^{\prime}_{1/2}z^{1/2}\bigr\rangle_{\infty}+\bigl\langle\mu^{\prime}z^{2}+\beta^{\prime}_{1}z\bigr\rangle_{\infty},$
with $\lambda^{\prime}\neq\mu^{\prime}$ and $\alpha^{\prime}_{3/2}\neq 0$.
The generic class of representations has thus rank 3, $k=2$ Fourier sphere coefficients, and the corresponding partition of the set of Stokes circles of $\bm{\breve{\Theta}}^{\rm II}_{\rm gen}$ is given by
@@
$N_{1}=N_{1}^{+}=\big\{\bigl\langle\lambda^{\prime}z^{2}+\alpha^{\prime}_{3/2}z^{3/2}+\alpha^{\prime}_{1}z+\alpha^{\prime}_{1/2}z^{1/2}\bigr\rangle_{\infty}\big\},$
$N_{2}=N_{2}^{-}=\big\{\bigl\langle\mu^{\prime}z^{2}+\beta^{\prime}_{1}z\bigr\rangle_{\infty}\big\}.$
4.3.3 Nongeneric representations
@@
The two principal subtrees of $\mathbf{T}^{\rm II}$ are not isomorphic, so the two nongeneric classes of representations have distinct forests. The standard representation corresponds to the nongeneric class of representations for $i=1$.
The second nongeneric class of representations has the following forest $\mathbf{F}^{\rm II}_{2}$:
113/23/21/21/22
It is of rank 2, with one irregular singularity (with Poincaré–Katz rank $3/2$) and one regular singularity.
Explicitly, elements of the orbit $\operatorname{SL}_{2}(\mathbb{C})\cdot\bm{\breve{\Theta}}^{\rm II}_{\rm gen}$ belonging to this class of representations are of the form
$\bm{\breve{\Theta}}^{\rm II}_{2}=\bigl\langle\lambda^{\prime\prime}z^{2}+\alpha^{\prime\prime}_{3/2}z^{3/2}+\alpha^{\prime\prime}_{1}z+\alpha^{\prime\prime}_{1/2}z^{1/2}\bigr\rangle_{\infty}+\bigl\langle 0\bigr\rangle_{a^{\prime\prime}},$
with $\alpha^{\prime\prime}\neq 0$, and $a^{\prime\prime}\in\mathbb{P}^{1}\smallsetminus\{\infty\}$, hence the corresponding non-modified irregular class is
$\bm{\Theta}^{\rm II}_{2}=\bigl\langle\lambda^{\prime\prime}z^{2}+\alpha^{\prime\prime}_{3/2}z^{3/2}+\alpha^{\prime\prime}_{1}z+\alpha^{\prime\prime}_{1/2}z^{1/2}\bigr\rangle_{\infty}+2\bigl\langle 0\bigr\rangle_{a^{\prime\prime}}.$
We notice that it corresponds precisely the other known Lax representation for Painlevé II due to Flaschka–Newell [26], also sometimes referred to as degenerate Painlevé IV.
4.3.4 Diagram
The corresponding diagram $\Gamma^{\rm II}$ is the affine $A_{1}$ Dynkin diagram [20, Section 7]
@@ 4.3.4 Diagram
The classes of representations correspond to different readings of the diagram, as indicated on the figure below
$\bigl\langle\alpha_{3}z^{3}+\cdots\bigr\rangle_{\infty}$$\bigl\langle\beta_{2}z^{2}+\cdots\bigr\rangle_{\infty}$$\bm{\breve{\Theta}}^{\rm II}_{\rm std}$$\bigl\langle\lambda^{\prime}z^{2}+\alpha^{\prime}_{3/2}z^{3/2}+\cdots\bigr\rangle_{\infty}$$\bigl\langle\mu^{\prime}z^{2}+\beta^{\prime}_{1}z\bigr\rangle_{\infty}$$\bm{\breve{\Theta}}^{\rm II}_{\rm gen}$$\bigl\langle\lambda^{\prime\prime}z^{2}+\alpha^{\prime\prime}_{3/2}z^{3/2}+\cdots\bigr\rangle_{\infty}$$\bigl\langle 0\bigr\rangle_{a^{\prime\prime}}$$\bm{\breve{\Theta}}^{\rm II}_{2}$
4.4 Painlevé III
4.4.1 Standard representation
The standard rank 2 Painlevé III Lax representation has two irregular singularities, both corresponding to a second order pole.
The corresponding fission forest $\mathbf{F}_{\rm std}^{\rm III(2)}$ is the following: [^1] We will denote Painlevé III, degenerate Painlevé III and doubly degenerate Painlevé III by $\rm III(2)$, $\rm III(1)$ and $\rm III(0)$, respectively, the numbers 2, 1, 0 referring to the number of parameters of each equation.
11
The irregular classes with this fission forest and minimal Poincaré–Katz ranks (both equal to 1) are the irregular classes of the form
$\bm{\Theta}^{\rm III(2)}_{\rm std}=\bigl\langle\alpha z_{b}^{-1}\bigr\rangle_{b}+\bigl\langle\beta z_{b}^{-1}\bigr\rangle_{b}+\bigl\langle\gamma z^{-1}_{a}\bigr\rangle_{a}+\bigl\langle\delta z_{a}^{-1}\bigr\rangle_{a},$
with $\alpha,\beta,\gamma,\delta\in\mathbb{C}$, $a,b\in\mathbb{P}^{1}$ distinct, $\alpha\neq\beta$, and $\gamma\neq\delta$.
If $(E,\nabla)$ is a connection on $\mathbb{P}^{1}$ of this type, the rank of its generic class of nearby representations is minimal when one of the singularities, say $b$, is at infinity, and, at the other singularity $a$ at finite distance, and one of the Stokes circles at finite distance, say $\bigl\langle\delta z_{a}^{-1}\bigr\rangle_{a}$, is the tame circle, with trivial formal monodromy (we can always reduce to this case by applying a Möbius transformation to send $b$ to infinity, and a twist by a rank one connection to set $\delta$ to zero and make the formal monodromy of $\bigl\langle\delta z_{a}^{-1}\bigr\rangle_{a}$ trivial).
This corresponds to having a modified irregular class of the form
$\bm{\breve{\Theta}}^{\rm III(2)}_{\rm std}=\bigl\langle\alpha z\bigr\rangle_{\infty}+\bigl\langle\beta z\bigr\rangle_{\infty}+\bigl\langle\gamma z_{a}^{-1}\bigr\rangle_{a}.$
with $\alpha\neq\beta$, and $\gamma\neq 0$. Let us determine the classes of nearby representations of $\bm{\breve{\Theta}}^{\rm III(2)}_{\rm std}$.
4.4.2 Generic representation
The short fission tree $\mathbf{T}^{\rm III(2)}$ of $\bm{\breve{\Theta}}^{\rm III(2)}_{\rm std}$ is the following:
11221/21/2
An irregular class of generic form has fission tree $\mathbf{T}^{\rm III(2)}$ if and only if it is of the form
$\bm{\breve{\Theta}}^{\rm III(2)}_{\rm gen}=\bigl\langle\lambda^{\prime}z^{2}+\alpha^{\prime}z\bigr\rangle_{\infty}+\bigl\langle\lambda^{\prime}z^{2}+\beta^{\prime}z\bigr\rangle_{\infty}+\bigl\langle\mu^{\prime}z^{2}+\gamma^{\prime}_{1}z+\gamma^{\prime}_{1/2}z^{1/2}\bigr\rangle_{\infty},$
with $\lambda^{\prime}\neq\mu^{\prime}$, $\alpha^{\prime}\neq\beta^{\prime}$, and $\gamma^{\prime}_{1/2}\neq 0$.
The generic class of representations has thus rank 4, $k=2$ Fourier sphere coefficients, and the corresponding partition of the set of Stokes circles of $\bm{\breve{\Theta}}^{\rm III(2)}_{\rm gen}$ is given by
$N_{1}=N_{1}^{-}=\big\{\bigl\langle\lambda^{\prime}z^{2}+\alpha^{\prime}z\bigr\rangle_{\infty},\bigl\langle\lambda^{\prime}z^{2}+\beta^{\prime}z\bigr\rangle_{\infty}\big\},$
$N_{2}=N_{2}^{-}=\big\{\bigl\langle\mu^{\prime}z^{2}+\gamma^{\prime}_{1}z+\gamma^{\prime}_{1/2}z^{1/2}\bigr\rangle_{\infty}\big\}.$
4.4.3 Nongeneric representations
Since the two principal subtrees are not isomorphic, the forests of the two corresponding nongeneric classes of representations do not coincide. The standard Lax representation is the nongeneric class of representations for $i=2$.
The other nongeneric class of representations, corresponding to $i=1$, has the following forest $\mathbf{F}^{\rm III(2)}_{1}$, with one irregular singularity (with Poincaré–Katz rank 1/2), and two regular singularities:
$1/2$
Explicitly, elements of the orbit $\operatorname{SL}_{2}(\mathbb{C})\cdot\bm{\breve{\Theta}}^{\rm III(2)}_{\rm gen}$ belonging to this class of representations are of the form
$\bm{\breve{\Theta}}^{\rm III(2)}_{1}=\bigl\langle 0\bigr\rangle_{a^{\prime\prime}}+\bigl\langle 0\bigr\rangle_{b^{\prime\prime}}+\bigl\langle\alpha^{\prime\prime}z^{1/2}\bigr\rangle_{\infty}$
with $a^{\prime\prime},b^{\prime\prime}\in\mathbb{P}^{1}\smallsetminus\{\infty\}$, $a^{\prime\prime}\neq b^{\prime\prime}$, $\alpha^{\prime\prime}\neq 0$, hence the corresponding non-modified irregular class is
$\bm{\Theta}^{\rm III(2)}_{1}=2\bigl\langle 0\bigr\rangle_{a^{\prime\prime}}+2\bigl\langle 0\bigr\rangle_{b^{\prime\prime}}+\bigl\langle\alpha^{\prime\prime}z^{1/2}\bigr\rangle_{\infty}.$
Notice that this class of representations corresponds to the alternative Lax representation for Painlevé III sometimes referred to as degenerate Painlevé V, cf. [36].
Remark 4.3.
We obtain in this way Figure 2 of the introduction: the fission forest on the left of the figure is $\mathbf{F}_{1}^{\rm III(2)}$, and the fission forest on the right of the figure is the forest associated to the modified irregular class $\bm{\breve{\Theta}}^{\rm III(2)}_{\rm std}$, i.e., compared to $\mathbf{F}_{\rm std}^{\rm III(2)}$ one full branch has been removed.
4.4.4 Diagram
The corresponding diagram $\Gamma^{\rm III(2)}$ is the following [20, Section 7]:
The classes of representations correspond to different readings, as represented on the figure below
$\bigl\langle\lambda^{\prime}z^{2}+\alpha^{\prime}z\bigr\rangle_{\infty}$$\bigl\langle\mu^{\prime}z^{2}+\gamma^{\prime}_{1}z+\gamma^{\prime}_{1/2}z^{1/2}\bigr\rangle_{\infty}$$\bigl\langle\lambda^{\prime}z^{2}+\beta^{\prime}z\bigr\rangle_{\infty}$$\bm{\breve{\Theta}}^{\rm III(2)}_{\rm gen}$$\bigl\langle\alpha z\bigr\rangle_{\infty}$$\bigl\langle\gamma z_{a}^{-1}\bigr\rangle_{a}$$\bigl\langle\beta z\bigr\rangle_{\infty}$$\bm{\breve{\Theta}}^{\rm III(2)}_{\rm std}$$\bigl\langle 0\bigr\rangle_{a^{\prime\prime}}$$\bigl\langle\alpha^{\prime\prime}z^{1/2}\bigr\rangle_{\infty}$$\bigl\langle 0\bigr\rangle_{b^{\prime\prime}}$$\bm{\breve{\Theta}}^{\rm III(2)}_{1}$
4.5 Degenerate Painlevé III
4.5.1 Standard representation
The standard Lax representation for degenerate Painlevé III, as for nondegenerate Painlevé III, has two irregular singularities, both corresponding to poles of order 2, but one of them with a nilpotent leading term giving rise in the Turritin–Levelt normal form to a single Stokes circle of slope $1/2$ [45].
This corresponds to the following fission forest $\mathbf{F}^{\rm III(1)}_{\rm std}$:
111/21/2
The irregular classes with fission forest $\mathbf{F}^{\rm III(1)}_{\rm std}$ and minimal Poincaré–Katz ranks (equal to 1 and 1/2) are those of the form
$\bm{\Theta}^{\rm III(1)}_{\rm std}=\bigl\langle\alpha z_{a}^{-1}\bigr\rangle_{a}+\bigl\langle\beta z_{a}^{-1}\bigr\rangle_{a}+\bigl\langle\gamma z_{b}^{-1/2}\bigr\rangle_{b}$
with $\in\mathbb{P}^{1}, aba\neq b$, $\gamma\neq 0$, and $\alpha\neq\beta$.
Given a connection of this type, the rank of its generic class of nearby representations is minimal if $b=\infty$ and one of the Stokes circles at $a$, say $\bigl\langle\beta z_{a}^{-1}\bigr\rangle_{a}$ is the tame circle, with trivial formal monodromy (we can always reduce to this case by applying a Möbius transformation to have $b=\infty$, and a twist by a rank one connection to set $\beta=0$ and make the formal monodromy of $\bigl\langle\beta z_{a}^{-1}\bigr\rangle_{a}$ trivial). This corresponds to having a modified irregular class of the form
$\bm{\breve{\Theta}}^{\rm III(1)}_{\rm std}=\bigl\langle\alpha z_{a}^{-1}\bigr\rangle_{a}+\bigl\langle\gamma z^{1/2}\bigr\rangle_{\infty}$
with $\alpha\neq 0$, $\gamma\neq 0$, $a\neq\infty$. Let us thus determine the classes of nearby representations of $\bm{\breve{\Theta}}^{\rm III(1)}_{\rm std}$.
4.5.2 Generic representation
The short fission tree $\mathbf{T}^{\rm III(1)}$ of $\bm{\breve{\Theta}}^{\rm III(1)}_{\rm std}$ is the following:
11221/21/2
An irregular class of generic form has fission tree $\mathbf{T}^{\rm III(1)}$ if and only if it is of the form
$\bm{\breve{\Theta}}^{\rm III(1)}_{\rm gen}=\bigl\langle\lambda^{\prime}z^{2}+\alpha^{\prime}_{1}z+\alpha^{\prime}_{1/2}z^{1/2}\bigr\rangle_{\infty}+\bigl\langle\mu^{\prime}z^{2}+\beta^{\prime}_{1}z+\beta^{\prime}_{1/2}z^{1/2}\bigr\rangle_{\infty}$
with $\lambda^{\prime}\neq\mu^{\prime}$, $\alpha^{\prime}_{1/2}\neq 0$ and $\beta^{\prime}_{1/2}\neq 0$.
The generic class of representations has thus rank 4, $k=2$ Fourier sphere coefficients, and the corresponding partition of the set of Stokes circles of $\bm{\breve{\Theta}}^{\rm III(1)}_{\rm gen}$ is given by
$N_{1}=N_{1}^{-}=\big\{\bigl\langle\lambda^{\prime}z^{2}+\alpha^{\prime}_{1}z+\alpha^{\prime}_{1/2}z^{1/2}\bigr\rangle_{\infty}\big\},$
$N_{2}=N_{2}^{-}=\big\{\bigl\langle\mu^{\prime}z^{2}+\beta^{\prime}_{1}z+\beta^{\prime}_{1/2}z^{1/2}\bigr\rangle_{\infty}\big\}.$
Furthermore, the two principal subtrees are isomorphic, so the two nongeneric classes of representations have the same forest, and there are only two distinct forests, the generic one and the nongeneric one, which coincides with the standard Lax representation.
4.5.3 Diagram
The corresponding diagram $\Gamma^{\rm III(1)}$ is the following [20, Section 7]:
441-11-1
The classes of representations correspond to different readings of the diagram, as indicated on the figure below